Calculate the mass of carbon in 10.00 g of glucose (C6H1206). Show the step by step process.

1
Expert's answer
2022-02-18T14:51:04-0500

Solution:

1) Calculate the molar mass of glucose (C6H1206):

M(C6H1206) = 6×Ar(C) + 12×Ar(H) + 6×Ar(O)

M(C6H1206) = 6×12.011 + 12×1.008 + 6×15.999 = 180.156 (g/mol)

M(C6H1206) = 180.156 g/mol


2) Calculate the mass percent of carbon (C) in glucose (C6H1206):

Mass percent of C = [6×Ar(C) / M(C6H1206)] × 100%

Mass percent of C = (6×12.011 / 180.156) × 100% = 40.00%


3) Calculate the the mass of carbon (C) in glucose (C6H1206):

Mass of C = (Mass percent of C × Mass of glucose) / 100%

Mass of C = (40.00% × 10.00 g) / 100% = 4 g

Mass of C = 4 g


Answer: The mass of carbon (C) is 4 grams